OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x2 + 6x + 7 = 0

OpenStudy (anonymous):

I got to :\[\frac{ 6 \pm \sqrt{8} }{ 2 }\]

hartnn (hartnn):

wouldn't it be -6 ? instead of 6 ?

OpenStudy (anonymous):

the choices are:\[36 \pm \sqrt{2}\] \[-6 \pm \sqrt{11}\] \[3 \pm \sqrt{11}\] \[-3 \pm \sqrt{2}\]

OpenStudy (anonymous):

yes sorry forgot the -6

hartnn (hartnn):

because u can simplify that, u know how ?

OpenStudy (anonymous):

no I don't know how to simplify it any further. I need to get to one of the above choices but don't understand how.

hartnn (hartnn):

\(\sqrt 8\) can be written as ?

OpenStudy (anonymous):

idk.. Where does the two go in the equation. I tried to divide it but 8/2 doesn't equal 2 or 11

hartnn (hartnn):

ok, maybe this will help. how can u factorize 8 ? \(\sqrt{a*a*a}=a\sqrt a\)

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

so u got the answer ?

OpenStudy (anonymous):

yes! it's \[-3 \pm \sqrt{2}\]

hartnn (hartnn):

thats correct :)